Solve the system by elimination. 6x - y = -5 & 4x - 2y = 6
Found 2 solutions by Alan3354, TimothyLamb:
Answer by Alan3354(69443)

Solve the system by elimination. 6x - y = -5 & 4x - 2y = 6
-------------
6x - y = -5
4x - 2y = 6 --> 2x - y = 3
--------
6x - y = -5
2x - y = 3
-------------------- Subtract

Answer by TimothyLamb(4379)

6x - y = -5
4x - 2y = 6
---
put the system of linear equations into standard form
---
6x - y = -5
4x - 2y = 6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -2
y = -7
---
now by elimination:
---
-2(6x - y = -5)
4x - 2y = 6
---
-12x + 2y = 10
4x - 2y = 6
---
add the two equations:
---
-8x = 16
x = -2
---
4x - 2y = 6
4(-2) - 2y = 6
-8 - 2y = 6
-2y = 14
y = -7
---
solution by elimination:
x = -2
y = -7
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---